# Distinguish dark matter theories with the cosmic web and next-generation surveys I: an alternative theory of gravity

### v2

@@
Table 1: Connectivity in the literature: From left to right, the columns give: data type, the size of the box, the resolution and the cosmic web tracers for simulations, the redshift of observed data, the persistence threshold to detect filaments, and the calibration method to set this free parameter of the DisPerSE code. If the persistence is not specified, it means that the skeleton has not been extracted with the DisPerSE code. One of the calibration methods consists of comparing the positions of peaks (CPmax) to those of the most massive halos such as galaxy groups and clusters ($>10^{13}\rm M_{\odot}$) (see section 2.4). Concerning cosmological simulations, all the cosmic web tracers are 3D quantities except for HzAGN simulation. For COSMOS and CFHTLS data, the calibration method involves optimally recovering the filaments in the 2D light-cone catalogues created from 3D cosmological simulations. 1 It corresponds to an absolute persistence cut because DisPerSE is not applied to discrete particles but on a 3d grid for the gas. These data are not taken into account in our study (for the reasons, see 2.5).
@@ 1 Introduction
Rather than modifying the nature of dark matter, one could question whether Poisson’s gravity provides a complete description of the gravity that dark matter undergoes on a cosmological scale. Even if the Poisson equation describes the large-scale structures of the Universe and agrees remarkably well with observations, alternatives to Poisson’s gravity can be considered to explain the $\Lambda$CDM problems. Here, we focus on a specific alternative theory of gravity, Monge-Ampère gravity. Replacing Poisson’s gravity with Monge-Ampère’s simply involves changing the trace operation in the Poisson equation to a determinant operation as:
@@
In this paper, we propose an approach to quantify how an alternative model of gravity alters the properties of the cosmic network as measured through the connectivity, i.e. the number of cosmic filaments branching out a cluster  [^1] Actually, we measure multiplicity and not connectivity, in the wording of Codis et al. (2018).. Connectivity is a tracer of the growth of structures and of the percolation properties of the cosmic web (Pichon et al. 2010) and as such its evolution throughout cosmic times is expected to informs us both about the equation of state of dark energy and the model of gravity which drives the coalescence of structures (Codis et al. 2018). Connectivity has recently emerged as an interesting probe of the large-scale structure around galaxies, groups and clusters, and was explored in several observational and numerical analyses. Pinpointing clusters, the most massive structures of the Universe, will ultimately mitigate the degeneracy between baryonic physics and the model of gravity. However this requires to observe sufficiently large area to beat the statistical uncertainties, and to overcome the fact that connectivity might be dependent on the degree of relaxation of the clusters (e.g. Gouin et al. 2021). We expect that having a relatively large sample size will help minimizing the risk of being contaminated by systematic errors of such nature, or at least controlling their impact.
@@ 2.1 DM-only cosmological simulations
The analysis presented in this paper uses snapshots of DM only simulations with Poisson gravity and Monge-Ampère gravity, starting from the same $\Lambda$CDM initial conditions using the N-GENIC code (Angulo et al. 2012; Springel et al. 2005). Our Poisson gravity simulation refers to classic $\Lambda$CDM cosmological simulations. These cosmological simulations from $z=49$ to $z=0$ were developed in a first work about Monge-Ampère gravity in a cosmological context (see Lévy et al. (2024)). With a mass resolution of $m_{\mathrm{DM}}=8\times 10^{9}\rm M_{\odot}$ and a uniform periodic-boundary cube of $L=205$ Mpc/h $\sim 300$ Mpc on a side, these simulations enable us to study the clustering of DM at different scales, from galaxies to filaments. The cosmological parameters that we used are taken from Planck Collaboration et al. (2016) results, with the matter density $\Omega_{\rm m}=0.3089$, the cosmological constant $\Omega_{\rm\Lambda}=0.6911$, the Hubble constant $H_{0}=67.74$ km s-1 Mpc-1 and $\sigma_{8}$ = 0.8159. The numerical implementation of this alternative gravity theory for a cosmological box using the latest algorithms based on optimal transport-based theory is detailed in (Lévy et al. 2024). Besides, the derivation of Monge-Ampère gravity from statistical physic principles is given in Brenier (2015) and Lévy et al. (2024).
@@ 2.2 Extracting dark matter halos
We used the adaptaHOP structure finder (Aubert et al. 2004; Tweed et al. 2009) in the most massive submaxima mode (MSM) to identify the halos in our simulations. We used a local threshold of 178 times the average matter density, with the local density of individual particles calculated using the 20 nearest neighbours. The MSM method is designed to construct the halo structure tree (host halo along with its subhalos and nested subhalos), ensuring that the halo itself encompasses the most massive local maximum. We require a minimum of 100 particles per (sub)halo. Figure 1 shows the halo mass function in our simulations at $z=0$. The alternative theory of gravity strongly affects the halo formation over cosmic time. Monge-Ampère simulation exhibits a lack of DM bound structures, especially at high masses ($>10^{13}\rm M_{\odot}$). We emphasise that Monge-Ampère equation is invariant under a larger group of symmetries, i.e. SL(3) or unimodular affine symmetry, than the Poisson equation which is invariant only under SO(3) and does not support deformation or shearing transformations. The extra symmetry facilitates the formation of anisotropic structures such as filaments, but it also makes them more resistant to collapse, which reduces the formation of halos (see Figure 1). This result can also be visually inspected on Figure 2. Given the halo mass function induced by the Monge-Ampère gravity, it appears challenging to reconcile it with the observed cluster mass function. This further confirms that this theory is not viable, even if we have not resolved baryonic physics in our simulations with Monge-Ampère gravity. Indeed, the calculated halo mass function in Figure 1 is sufficient to exclude this gravity theory. Indeed, there is a difference of one to two orders of magnitude between the predictions of Poisson and Monge-Ampère. However, the observations with their error bars regarding the halo mass function are tightly constrained in our considered mass range and are in very good agreement with $\Lambda$CDM (Driver et al. 2022). On the contrary, our halo mass function in the Monge-Ampère gravity is clearly in disagreement with $\Lambda$CDM, considering the observation error bars. As mentioned before, our approach aims to find new metrics to constrain alternative gravity theories, as well as alternative dark matter theories that are consistent with certain observations, such as the stellar or cluster mass function.
@@ 2.5 Connectivity $\kappa$
By design, we are mostly concerned here by the connectivity of galaxy groups and clusters given the resolution of our simulations and the next-generation surveys. One of the objectives of this paper is to investigate whether the connectivity can be a good metric to validate or invalidate alternative gravity theories such as Monge-Ampère gravity in a cosmological context and ultimately extend it to alternative DM theories such as WDM, SIDM, and FDM. The local number of filaments that are connected to galaxy groups and clusters in a sphere of 1.5 rvir is the connectivity $\kappa$, similarly to Darragh Ford et al. (2019); Gouin et al. (2021); Sarron et al. (2019) [^2] Actually, we measure multiplicity and not connectivity, in the wording of Codis et al. (2018).. Given that the choice of a radius of 1.5 times the virial radius is arbitrary for calculating the mean connectivity, measurements within 1 and 2 times the virial radius are considered as error bars in all our figures.
@@
Figure 5 also compares the mean connectivity of Poisson and Monge-Ampère theories extracted from the DM particles. As the extra symmetry enhances the formation of filaments in Monge-Ampère gravity, it results that the mean connectivity is higher at high halo mass compared to Poisson gravity. We also stress that in both gravity theories, the connectivity increases with mass in agreement with previous studies. Besides, the trend of the mean connectivity approaching zero at low halo mass must be attributed to the lack of resolution in our simulations (see Figure 6). We emphasize that we take into account halos that have zero connectivity according to DisPerSE code. Figure 6 contains the entire set of observed data that we have compiled from the literature (see details in Table 1). In the case of the cosmic web extracted from dark matter halos, there is still a deviation between Poisson connectivity and that of Monge-Ampère at high halo mass (see Figure 2). This behaviour persists because a cosmic web estimated from halos slightly underestimates the actual connectivity revealed by the DM density field (see Figure 7). The connectivity based on halos as a function of connectivity based on DM particles was computed by fitting the mean connectivities presented in the Figures 5 and  6 by excluding the points from the zero-slope region. We also examined whether the connectivity of the same halos identified in both simulations was correlated. This analysis, applied to 742 halos at $z=0$, revealed a positive correlation, as expected from Figure 6, but it remains very weak between halos formed in Poisson gravity and Monge-Ampère gravity. The simulation data concerning the connectivity used in our study of Poisson and Monge-Ampère models are available here  [^4] https://www.iap.fr/useriap/boldrini/data.html.
@@ 3.2 Alternative DM theories and cosmic web
A question that now arises concerns the viability of our approach not only for gravity theories affecting the dynamics of matter, but for alternative theories of dark matter whose intrinsic properties are modified. Intuitively, one would expect our new metrics to work within the framework of DM theories that affect the cosmic web and, consequently, the connectivity of filaments. Indeed, we have highlighted that modifying the gravity in Monge-Ampère theory notably impacts large-scale dynamics, and this modification leaves an imprint on the connectivity. Are there alternative DM theories with exotic properties that affect the cosmic skeleton? A wide range of alternative DM models has been proposed over the last decades. Two main classes of alternative DM models have been simulated on large scales: warm dark matter (WDM, Bode et al. 2001) and fuzzy dark matter (FDM, Hu et al. 2000). These theories have gained attention as alternatives to the standard CDM model because they offer potential solutions to some of CDM puzzles, particularly on small scales (Bullock & Boylan-Kolchin 2017; Boldrini 2021). That’s why there are few studies, especially on large scales, concerning the cosmic web and its properties.
@@
Since particles in WDM and FDM models have different intrinsic properties from the CDM particle candidates, the effect of these particles on structure formation and evolution is expected to be qualitatively different on both small and large scales and particularly on the overdense structures that connect halos (Paduroiu et al. 2015; Mocz et al. 2020; May & Springel 2023). In WDM cosmology, it has been demonstrated that the presence of less substructures is associated with warmer dark matter. Specifically, as the particle velocity dispersion increases, the filaments become larger, and their preservation extends over a longer period. For WDM models, structure formation follows a complex scenario in which both top-down and bottom-up processes play a role (Paduroiu et al. 2015). Due to this hybrid structure formation, we expect that this difference will be evident in the connectivity across redshifts. Like WDM, FDM filaments do not fragment into halos like those of CDM (Mocz et al. 2020; May & Springel 2023). Over time, CDM filaments can fragment and collapse to form gravitationally bound structures known as DM halos, which are seeds for the formation of galaxies and galaxy clusters. This is why WDM and FDM simulations appear similarly filamentary (Mocz et al. 2020). Furthermore, these alternative models can also induce a delay in the formation of structures. In CDM scenario, the first structures begin to form well before reaching $z=9$. In contrast, in a WDM universe, the first collapse may be delayed until $z=4$ (Paduroiu et al. 2015). All these physical arguments from the literature invite us to conduct an investigation into the properties of filaments in these alternative theories, following a benchmark that we propose.

### v3

@@
Notes: From left to right, the columns give: data type, the size of the box, the resolution and the cosmic web tracers for simulations, the redshift of observed data, the persistence threshold to detect filaments, and the calibration method to set this free parameter of the DisPerSE code. If the persistence is not specified, it means that the skeleton has not been extracted with the DisPerSE code. One of the calibration methods consists of comparing the positions of peaks (CPmax) to those of the most massive halos such as galaxy groups and clusters ($>10^{13}\rm M_{\odot}$) (see section 2.4). Concerning cosmological simulations, all the cosmic web tracers are 3D quantities except for HzAGN simulation. For COSMOS and CFHTLS data, the calibration method involves optimally recovering the filaments in the 2D light-cone catalogues created from 3D cosmological simulations. 1 It corresponds to an absolute persistence cut because DisPerSE is not applied to discrete particles but on a 3d grid for the gas. These data are not taken into account in our study (for the reasons, see 2.5).
@@ 1 Introduction
Rather than modifying the nature of DM, one could question whether Poisson’s gravity provides a complete description of the gravity that DM undergoes on a cosmological scale. Even if the Poisson equation describes the large-scale structures of the Universe and agrees remarkably well with observations, alternatives to Poisson’s gravity can be considered to explain the $\Lambda$CDM problems. Here, we focus on a specific alternative theory of gravity, Monge-Ampère gravity. Replacing Poisson’s gravity with Monge-Ampère’s simply involves changing the trace operation in the Poisson equation to a determinant operation as:
@@
In this paper, we propose an approach to quantify how an alternative model of gravity alters the properties of the cosmic network as measured through the connectivity, i.e. the number of cosmic filaments branching out a cluster. Actually, we measure multiplicity and not connectivity, in the wording of Codis et al. (2018). Connectivity is a tracer of the growth of structures and of the percolation properties of the cosmic web (Pichon et al. 2010) and as such its evolution throughout cosmic times is expected to informs us both about the equation of state of dark energy and the model of gravity which drives the coalescence of structures (Codis et al. 2018). Connectivity has recently emerged as an interesting probe of the large-scale structure around galaxies, groups and clusters, and was explored in several observational and numerical analyses. Pinpointing clusters, the most massive structures of the Universe, will ultimately mitigate the degeneracy between baryonic physics and the model of gravity. However this requires to observe sufficiently large area to beat the statistical uncertainties, and to overcome the fact that connectivity might be dependent on the degree of relaxation of the clusters (e.g. Gouin et al. 2021). We expect that having a relatively large sample size will help minimizing the risk of being contaminated by systematic errors of such nature, or at least controlling their impact.
@@ 2.1 DM-only cosmological simulations
The analysis presented in this paper uses snapshots of DM only simulations with Poisson gravity and Monge-Ampère gravity, starting from the same $\Lambda$CDM initial conditions using the N-GENIC code (Angulo et al. 2012; Springel et al. 2005). Our Poisson gravity simulation refers to classic $\Lambda$CDM cosmological simulations. These cosmological simulations from $z=49$ to $z=0$ were developed in a first work about Monge-Ampère gravity in a cosmological context (see Lévy et al. (2024)). With a mass resolution of $m_{\mathrm{DM}}=8\times 10^{9}\rm M_{\odot}$ and a uniform periodic-boundary cube of $L=205$ Mpc/h $\sim 300$ Mpc on a side, these simulations enable us to study the clustering of DM at different scales, from galaxies to filaments. The cosmological parameters that we used are taken from Planck Collaboration et al. (2016) results, with the matter density $\Omega_{\rm m}=0.3089$, the cosmological constant $\Omega_{\rm\Lambda}=0.6911$, the Hubble constant $H_{0}=67.74$ km s-1 Mpc-1 and $\sigma_{8}$ = 0.8159. The numerical implementation of this alternative gravity theory for a cosmological box using the latest algorithms based on optimal transport-based theory is detailed in (Lévy et al. 2024). Besides, the derivation of Monge-Ampère gravity from statistical physic principles is given in Brenier (2015) and Lévy et al. (2024). At each time step, the Monge-Ampère equation is solved using the optimal transport algorithm (fully detailed in Lévy, Bruno (2015)). Then, gravitational forces are computed and momenta and positions are updated, as in standard N-body codes such as Gadget. Ultimately, the only difference from a classic N-body approach lies in the definition and calculation of the gravitational force. The large-scale structures have globally the same shape for both simulations, which was expected since Monge-Ampère theory can be considered as a small perturbation of Poisson gravity. As with Poisson gravity, MAG is a theory built without free parameters. This is why the two are identical in 1D. While one of its limitations is that it cannot be tuned, this also makes it a powerful framework for testing our approach. According to the power spectra (shown in Figure 9 of Lévy et al. 2024), at large scales, $\Lambda$CDM and this alternative gravity theory have the same trend since in low density regions far away from dense regions, the nonlinearities are negligable. However, on small scales, Monge-Ampère has much less power than $\Lambda$CDM. This behavior can also be observed through the percentage difference between Monge-Ampère and CDM at different scales shown in Figure 11.
@@ 2.2 Extracting dark matter halos
We used the adaptaHOP structure finder (Aubert et al. 2004; Tweed et al. 2009) in the most massive submaxima mode (MSM) to identify the halos in our simulations. We used a local threshold of 178 times the average matter density, with the local density of individual particles calculated using the 20 nearest neighbours. The MSM method is designed to construct the halo structure tree (host halo along with its subhalos and nested subhalos), ensuring that the halo itself encompasses the most massive local maximum. We require a minimum of 100 particles per (sub)halo. Figure 1 shows the halo mass function in our simulations at $z=0$. The alternative theory of gravity strongly affects the halo formation over cosmic time. Monge-Ampère simulation exhibits a lack of DM bound structures, especially at high masses ($>10^{13}\rm M_{\odot}$). We emphasise that Monge-Ampère equation is invariant under a larger group of symmetries, i.e. SL(3) or unimodular affine symmetry, than the Poisson equation which is invariant only under SO(3) and does not support deformation or shearing transformations. The extra symmetry facilitates the formation of anisotropic structures such as filaments, but it also makes them more resistant to collapse, which reduces the formation of halos (see Figure 1). This result can also be visually inspected on Figure 2. Given the halo mass function induced by the Monge-Ampère gravity, it appears challenging to reconcile it with the observed cluster mass function. This further confirms that this theory is not viable, even if we have not resolved baryonic physics in our simulations with Monge-Ampère gravity. Indeed, the calculated halo mass function in Figure 1 is sufficient to exclude this gravity theory. Indeed, there is a difference of one to two orders of magnitude between the predictions of Poisson and Monge-Ampère. However, the observations with their error bars regarding the halo mass function are tightly constrained in our considered mass range and are in very good agreement with $\Lambda$CDM (Driver et al. 2022). On the contrary, our halo mass function in the Monge-Ampère gravity is clearly in disagreement with $\Lambda$CDM, considering the observation error bars. As mentioned before, our approach aims to find new metrics to constrain alternative gravity theories, as well as alternative DM theories that are consistent with certain observations, such as the stellar or cluster mass function.
@@ 2.5 Connectivity $\kappa$
By design, we are mostly concerned here by the connectivity of galaxy groups and clusters given the resolution of our simulations and the next-generation surveys. One of the objectives of this paper is to investigate whether the connectivity can be a good metric to validate or invalidate alternative gravity theories such as Monge-Ampère gravity in a cosmological context and ultimately extend it to alternative DM theories such as WDM, SIDM, and FDM. The local number of filaments that are connected to galaxy groups and clusters in a sphere of 1.5 rvir is the connectivity $\kappa$, similarly to Darragh Ford et al. (2019); Gouin et al. (2021); Sarron et al. (2019). Given that the choice of a radius of 1.5 times the virial radius is arbitrary for calculating the mean connectivity, measurements within 1 and 2 times the virial radius are considered as error bars in all our figures.
@@
[…]
Figure 5 also compares the mean connectivity of Poisson and Monge-Ampère theories extracted from the DM particles. As the extra symmetry enhances the formation of filaments in Monge-Ampère gravity, it results that the mean connectivity is higher at high halo mass compared to Poisson gravity. We also stress that in both gravity theories, the connectivity increases with mass in agreement with previous studies. Besides, the trend of the mean connectivity approaching zero at low halo mass must be attributed to the lack of resolution in our simulations (see Figure 6). We emphasize that we take into account halos that have zero connectivity according to DisPerSE code. Figure 6 contains the entire set of observed data that we have compiled from the literature (see details in Table 1). In the case of the cosmic web extracted from DM halos, there is still a deviation between Poisson connectivity and that of Monge-Ampère at high halo mass (see Figure 2). This behaviour persists because a cosmic web estimated from halos slightly underestimates the actual connectivity revealed by the DM density field (see Figure 7). The connectivity based on halos as a function of connectivity based on DM particles was computed by fitting the mean connectivities presented in the Figures 5 and  6 by excluding the points from the zero-slope region. We also examined whether the connectivity of the same halos identified in both simulations was correlated. This analysis, applied to 742 halos at $z=0$, revealed a positive correlation, as expected from Figure 6, but it remains very weak between halos formed in Poisson gravity and Monge-Ampère gravity. The choice to measure connectivity on halos rather than on the density field is driven by the ability to apply the same methods to observable galaxy data, as previously done by Darragh Ford et al. (2019); Gouin et al. (2021); Kraljic et al. (2020); Galárraga-Espinosa et al. (2023a), and as currently implemented within the Euclid consortium. The simulation data concerning the connectivity used in our study of Poisson and Monge-Ampère models are available here  [^2] https://www.iap.fr/useriap/boldrini/data.html.
@@ 3.2 Alternative DM theories and cosmic web
A question that now arises concerns the viability of our approach not only for gravity theories affecting the dynamics of matter, but for alternative theories of DM whose intrinsic properties are modified. Intuitively, one would expect our new metrics to work within the framework of DM theories that affect the cosmic web and, consequently, the connectivity of filaments. Indeed, we have highlighted that modifying the gravity in Monge-Ampère theory notably impacts large-scale dynamics, and this modification leaves an imprint on the connectivity. Are there alternative DM theories with exotic properties that affect the cosmic skeleton? A wide range of alternative DM models has been proposed over the last decades. Two main classes of alternative DM models have been simulated on large scales: warm DM (WDM, Bode et al. 2001) and fuzzy DM (FDM, Hu et al. 2000). These theories have gained attention as alternatives to the standard CDM model because they offer potential solutions to some of CDM puzzles, particularly on small scales (Bullock & Boylan-Kolchin 2017; Boldrini 2021). That’s why there are few studies, especially on large scales, concerning the cosmic web and its properties.
@@
Since particles in WDM and FDM models have different intrinsic properties from the CDM particle candidates, the effect of these particles on structure formation and evolution is expected to be qualitatively different on both small and large scales and particularly on the overdense structures that connect halos (Paduroiu et al. 2015; Mocz et al. 2020; May & Springel 2023). In WDM cosmology, it has been demonstrated that the presence of less substructures is associated with warmer DM. Specifically, as the particle velocity dispersion increases, the filaments become larger, and their preservation extends over a longer period. For WDM models, structure formation follows a complex scenario in which both top-down and bottom-up processes play a role (Paduroiu et al. 2015). Due to this hybrid structure formation, we expect that this difference will be evident in the connectivity across redshifts. Like WDM, FDM filaments do not fragment into halos like those of CDM (Mocz et al. 2020; May & Springel 2023). Over time, CDM filaments can fragment and collapse to form gravitationally bound structures known as DM halos, which are seeds for the formation of galaxies and galaxy clusters. This is why WDM and FDM simulations appear similarly filamentary (Mocz et al. 2020). Furthermore, these alternative models can also induce a delay in the formation of structures. In CDM scenario, the first structures begin to form well before reaching $z=9$. In contrast, in a WDM universe, the first collapse may be delayed until $z=4$ (Paduroiu et al. 2015).
Figure 11 presents a comparison of power spectra normalized to CDM for the Monge-Ampère model and two alternative DM (ADM) theories as a function of the spatial scale of structures at $z=3$ and 0. Given that the deviations in the power spectrum of the Monge-Ampère model from CDM are sufficient to induce visible differences in halo connectivity, we use this to infer the impact of FDM and WDM spectra on connectivity by comparing their respective deviations from CDM. By visual inspection, the WDM spectrum at $z=0$ from Viel et al. (2012) also shows a suppression of DM clustering, which is more pronounced at small scales than in the Monge-Ampère model. Unlike WDM, where the cutoff is gradual, FDM at $z=3$ from May & Springel (2023) exhibits a sharp cutoff at small scales due to quantum pressure effects. Therefore, we expect these alternative DM theories to disrupt the connectivity of their halos, as they exhibit cutoffs similar to the gravity theory investigated here but at different scales (see Figure 11). Indeed, in WDM and FDM scenarios, a suppression of the number of halos is expected compared to the CDM model, particularly for $M<=10^{9}M_{\odot}$, which will disrupt the connectivity of more massive halos because connectivity is simply the count, within a given radius, of the number of filaments traced by the surrounding halos. All these physical arguments from the literature motivate us to investigate the properties of filaments in these alternative theories, under a benchmark framework that we propose. We emphasize that the expected impact on connectivity remains a hypothesis to be tested, which reinforces the need for further studies.
@@ 3.3 Forecasts for future surveys
Figure 11: Percentage difference between alternative DM theories (ADM) and CDM power spectra. The power spectra data for FDM and WDM are taken from May & Springel (2023) and Viel et al. (2012) with a DM mass parameter of 2.5 ×$\times 10^{-22}$ eV and 2 keV at $z=3$ and $z=0$, respectively. The cutoffs observed in WDM and FDM, similar to those in the Monge-Ampère model but occurring at different scales, are expected to induce differences in halo connectivity compared to CDM.